A rock falls from a tower that is 400 ft high. AS it is falling, its height is given by formula h = 400 - 16t^2. How many seconds will it takes for the rock to hit the ground (h = 0)? Show your work.
Answer by mananth(16949)

since h =0 when it touches the ground
400-16t^2=0
16t^2=400
t^2=400/16
t^2= 25
t=5 seconds